# MEGA-Bench[Uncaptioned image]: Scaling Multimodal Evaluation to over 500 Real-World Tasks

### v1

@@ 1 Introduction
Large foundation models (OpenAI, 2023; 2024a; Anthropic, 2024; Google, 2023; Meta, 2024; Alibaba, 2024) have dramatically transformed the landscape of artificial intelligence by showcasing exceptional capabilities across various tasks and domains. Originating in the realm of natural language processing, these models have progressively expanded to perceive and interpret multimodal information, including single images, multiple images, and videos. Previously, multimodal models were mainly used for standardized tasks like image captioning (Lin et al., 2014), video captioning (Wang et al., 2019), and visual question answering (Antol et al., 2015; Goyal et al., 2017; Xiao et al., 2021). With the recent progress on multimodal alignment, these models have shown great potential to solve any desired task with a well-designed prompt. As a result, people have applied them to assist with many realistic tasks like “web navigation” (Koh et al., 2024), “game playing” (Valevski et al., 2024), “travel planning” (Xie et al., 2024), “visual navigation” (Wang et al., 2023a), “sports analysis” (Xia et al., 2024), “visual entity recognition” (Hu et al., 2023), “visual quality assessment” (Ku et al., 2024), and more. These efforts have significantly increased the utility of multimodal models.
@@
1. Among flagship models, GPT-4o is the current best model on a wide spectrum of multi-modal tasks, with a significant margin (3.5%) over the second best (Claude 3.5).
@@
4. Proprietary models can effectively leverage Chain-of-Thought (CoT) prompting to improve their performance, while open-source models hardly produce helpful reasoning processes. In our evaluation results, 8 out of 11 open-source models get worse results with CoT prompting.
@@ 3.1 Benchmark Construction Process
We then distribute the second-level nodes in “draft” taxonomy tree to the annotators based on their preferences. This top-down annotation framework can help minimize the overlaps of tasks contributed by different annotators. To ensure consistent commitment throughout the project, we call up over 16 designated expert annotators. Our annotators are all graduate students or above from different majors like Computer Science, Electronics, Bio-statistics, and Finance. The annotators can 1) refine the “draft” taxonomy by adding/deleting nodes, 2) add “task group” nodes and then add a series of tasks under that, and 3) directly add tasks under the second-level meta-task node. To simplify the task contribution, we develop tools to facilitate the annotation process, including 1) an interactive annotation tool that defines the annotation format and automatically unifies all annotations as JSON files (Please see §A.1 for our annotation format), 2) a GitHub repository to coordinate the task submission, reviewing, and discussion process, which was inspired by BIG-bench (Srivastava et al., 2022), and 3) a visualization tool that allows annotators to browse the existing tasks and the evaluation results of representative vision language models (VLMs). We coordinate all the annotators to ensure they understand our expectations and continuously improve our tools.
@@ 3.2 Metrics for answers in diverse output formats
Rule-based metrics. When there is a unique answer under the question context or the correctness of the answer can be verified by rules (e.g., if the generated story/poetry meets the desired formats, or if the generated code can pass the given test cases), we implement rule-based metrics for evaluation. To satisfy the needs of all tasks submitted by our annotators, we end up with a suite of over 40 rule-based metrics. Robust string parsing is also implemented to extract the answer from the model’s response. We conduct a sanity check to ensure the correct implementation of all rule-based metrics. Specifically, we create an “oracle” model that always returns the ground truth annotation. We then compute its score over all tasks evaluated with rule-based metrics. The sanity check is passed when the “oracle” model gets a 1.0 score. See §C.4 for details.
LLM-assisted metrics. For open-ended tasks that do not have a unique answer, we instead employ an LLM-assisted metric (Zheng et al., 2023; Li et al., 2023a). We design a per-task evaluation prompt template and fill in the tailored evaluation criteria for each task. The LLM is instructed to compare the model response with the reference answer and assign a score from 1 to 10. The score is then normalized into [0, 1] to be consistent with the other metrics. See §C.3 for details.
@@ 3.3 Multi-dimensional Keywords for fine-grained analysis
MVBench (Li et al., 2024d) Repurposed Existing Video Multi-choice (MC) 1 20
@@ 3.4 Dataset statistics and comparison with other benchmarks
MEGA-Bench contains 505 real-world tasks with 8,186 manually annotated or repurposed samples. Even for repurposed data, considerable effort is needed to convert the original annotations into specific task descriptions, diverse output formats, and additional instructions to include auxiliary information and about the formatting. Figure 2 shows the task distribution of all five dimensions, and the detailed breakdown statistics are in §B.2.
@@ 4 Experiments
We evaluate 16 VLMs with multi-image support on MEGA-Bench. §4.1 describes the models being assessed and the evaluation pipeline. §4.2 presents the quantitative evaluation results with a fine-grained analytical breakdown. §4.3 discusses the impact of the number of examples per task on the evaluation results. We provide detailed inspections of model behaviors with typical error cases in §E of the Appendix.
@@ 4.1 Evaluation Settings
Evaluated models. We evaluate a diverse range of large multimodal models. The proprietary models assessed include GPT-4o (0513) and GPT-4o mini (OpenAI, 2024a), Claude-3.5-Sonnet (Anthropic, 2024), Gemini-1.5-Pro and Gemini-1.5-Flash (002) (Google, 2024a). For open-source models, we focus on large flagship ($>$70B parameters) and medium-scale (7B-15B parameters) models. The large models include Qwen2-VL-72B (Alibaba, 2024), InternVL2-Llama3-76B (Chen et al., 2024d), and LLaVA-OneVision-72B (Li et al., 2024a), while medium-scale models comprise Qwen2-VL-7B (Alibaba, 2024), Pixtral 12B (Mistral, 2024), InternVL2-8B (Chen et al., 2024d), Phi-3.5-Vision (Abdin et al., 2024), MiniCPM-V2.6 (Yao et al., 2024), LLaVA-OneVision-7B (Li et al., 2024a), Llama-3.2-11B Meta (2024), and Idefics3-8B-Llama3 Laurençon et al. (2024).
Evaluation pipeline. MEGA-Bench has diverse and flexible formats. To ensure the models have clear instructions on the output format, we provide all evaluated VLMs with a one-shot in-context example. For each query, we fill in a pre-defined prompt template with the task instructions written by the task annotators, the 1-shot example, and the concrete query question. Since this one-shot example’s primary purpose is to illustrate the output format, we allocate only a tiny portion of the total image budget for it. For each model, we conduct experiments with and without Chain-of-Thought (CoT) prompting (Wei et al., 2022) for the Core tasks (the one-shot example of Open-ended tasks already contains CoT demonstrations). The prompt templates and other evaluation details (e.g., the frame sampling strategy for video inputs) are in §C.
Table 2: The main results of different models on the Core and Open-ended subset of MEGA-Bench, with 440 and 65 tasks, respectively. We report the macro mean scores across all tasks in each set. The overall score is the weighted average of the Core and Open-ended scores. When computing the overall score, we use the higher Core score from ‘w/o CoT’ and ‘w/ CoT’.
Model Eval Tier Open Source Core (rule-based eval) Open-ended (GPT eval) Overall
@@
GPT-4o (0513) (OpenAI, 2024a) Flagship No 51.88 52.52 64.78 54.10
Claude-3.5-Sonnet (Anthropic, 2024) Flagship No 48.63 50.24 63.74 51.97
Gemini-1.5-Pro-002 (Google, 2024b) Flagship No 46.89 48.14 58.58 49.48
Gemini-1.5-Flash-002 (Google, 2024b) Efficiency No 41.84 41.84 56.91 43.78
GPT-4o mini (OpenAI, 2024b) Efficiency No 39.74 40.71 58.65 43.02
Qwen2-VL-72B (Alibaba, 2024) Flagship Yes 46.24 45.28 56.40 47.55
InternVL2-Llama3-76B (Chen et al., 2024d) Flagship Yes 34.98 35.54 51.93 37.65
LLaVA-OneVision-72B (Li et al., 2024a) Flagship Yes 31.96 29.73 45.99 33.77
Qwen2-VL-7B (Alibaba, 2024) Efficiency Yes 34.73 32.84 43.96 35.91
Pixtral 12B (Mistral, 2024) Efficiency Yes 31.87 31.32 45.66 33.64
InternVL2-8B (Chen et al., 2024d) Efficiency Yes 25.92 24.06 39.79 27.71
Phi-3.5-Vision (Abdin et al., 2024) Efficiency Yes 23.24 22.95 39.48 25.33
MiniCPM-V2.6 (Yao et al., 2024) Efficiency Yes 22.84 22.90 41.73 25.32
LLaVA-OneVision-7B (Li et al., 2024a) Efficiency Yes 22.39 21.35 33.98 23.88
Llama-3.2-11B (Meta, 2024) Efficiency Yes 10.04 15.98 31.73 18.01
@@ 4.2 Main Results with breakdown analysis
Flagship models. Unlike the results on recent benchmarks like MMMU-Pro (Yue et al., 2024b) where GPT-4o and Claude-3.5 get close scores, GPT-4o outperforms Claude-3.5 with a clear margin on MEGA-Bench. Investigating the breakdown results, we observe that GPT-4o wins in most applications/skills, except for coding, math, and planning-related tasks, where the answers are typically in a “structured” output format, as described in Figure 5. One typical behavior of Claude-3.5 we observed is that it frequently refuses to answer routine knowledge or commonsense questions, such as the name and nationality of famous actors. The bottom radar maps show that Claude-3.5 has the best ethical/safety reasoning score but falls behind in knowledge and information extraction, partially because it refuses to answer many queries. Our Planning application keyword contains tasks related to symbolic planning (Zhu et al., 2024), navigation (Ku et al., 2020), chess games (Fu et al., 2024b), puzzle games (e.g., maze, Sudoku, etc.), and even the best models get low scores.
@@
Efficiency models. Figure 6 analyzes the results on efficiency models. In general, Gemini-1.5-flash-002 has the best performance with exceptional scores in Science and Metrics applications. The Metrics keyword contains tasks such as rating the quality of GenAI results (He et al., 2024; Jiang et al., 2024b) and requires deep multimodal reasoning and commonsense.
Chain-of-Thought. An interesting finding is that the CoT prompt (See §C) effectively guides all proprietary models to generate a detailed reasoning process, and flagship-tier proprietary models all obtain better performance on the Core set. However, it has almost no effect on most open-source models. For example, the Qwen2-VL, InternVL2, and LLaVA-OneVision models rarely produce reasoning when given a CoT instruction, and sometimes get confused about the required format after generating the reasoning process, leading to a lower score on the Core set.
Some open-source models get comparatively low scores for their parameter count. Llama-3.2-11B has difficulty leveraging the one-shot example to understand the correct output format and tends to generate a long descriptive sentence instead. This issue is alleviated under the CoT setting because the prompt provides extra instructions on the output format beyond the one-shot example, requesting the model to strictly separate the reasoning process from the final answer. Idefics3 frequently repeats the example answer from the one-shot demonstration. We suspect the reason for this problem is the poor support for multi-image (as our query contains at least two images, including the one-shot example) since it can generate reasonable responses with a single-image input.
@@ 4.3 Analysis on the number of samples per task
Figure 7: (Left) The bootstrap distribution of benchmark scores obtained by a subset of models, as we gradually increase the bootstrap sample size of the number of examples per task. (Right) The task-wise error distribution of GPT-4o (0513) over a subset of 255 Core tasks.
As discussed in §1, one of MEGA-Bench’s goals is to optimize the inference cost. We prioritize expanding the number of tasks over adding many examples per task. To understand the robustness of the benchmark score with around 15 examples per task, we obtained bootstrap distributions (Efron & Tibshirani, 1994; Hesterberg, 2011) of the model scores for our Core set without CoT prompting. We did this by taking a random subset of the model’s responses of size $n$ with replacement for each task and calculating the macro-mean scores. To ensure that the bootstrap distribution was numerically stable, we ran 10,000 Monte Carlo simulations. Figure 7 (left) shows that the variance in model scores rapidly narrows as the number of examples per task increases. As the number of examples per task increases beyond 7, the marginal return in variance reduction diminishes.

### v2

@@ 1 Introduction
Large foundation models (OpenAI, 2023; 2024a; Anthropic, 2024a; Google, 2023; Meta, 2024; Alibaba, 2024) have dramatically transformed the landscape of artificial intelligence by showcasing exceptional capabilities across various tasks and domains. Originating in the realm of natural language processing, these models have progressively expanded to perceive and interpret multimodal information, including single images, multiple images, and videos. Previously, multimodal models were mainly used for standardized tasks like image captioning (Lin et al., 2014), video captioning (Wang et al., 2019), and visual question answering (Antol et al., 2015; Goyal et al., 2017; Xiao et al., 2021). With the recent progress on multimodal alignment, these models have shown great potential to solve any desired task with a well-designed prompt. As a result, people have applied them to assist with many realistic tasks like “web navigation” (Koh et al., 2024), “game playing” (Valevski et al., 2024), “travel planning” (Xie et al., 2024), “visual navigation” (Wang et al., 2023a), “sports analysis” (Xia et al., 2024), “visual entity recognition” (Hu et al., 2023), “visual quality assessment” (Ku et al., 2024), and more. These efforts have significantly increased the utility of multimodal models.
@@
1. Among flagship models, Claude 3.5 Sonnet (1022) and GPT-4o (0513) currently lead in performance across a wide range of multimodal tasks, with less than a 0.1% difference in their overall scores. Detailed breakdown analyses show Claude 3.5 Sonnet excels in planning and math, while GPT-4o leads in information extraction and knowledge-intensive tasks.
@@
4. Proprietary models can effectively leverage Chain-of-Thought (CoT) prompting to improve their performance, while open-source models hardly produce helpful reasoning processes. In our evaluation results, 11 out of 16 open-source models get worse results with CoT prompting.
@@ 3.1 Benchmark Construction Process
We then distribute the second-level nodes in “draft” taxonomy tree to the annotators based on their preferences. This top-down annotation framework can help minimize the overlaps of tasks contributed by different annotators. To ensure consistent commitment throughout the project, we call up over 16 designated expert annotators. Our annotators are all graduate students or above from different majors like Computer Science, Electronics, Bio-statistics, and Finance. The annotators can 1) refine the “draft” taxonomy by adding/deleting nodes, 2) add “task group” nodes and then add a series of tasks under that, and 3) directly add tasks under the second-level meta-task node. To simplify the task contribution, we develop tools to facilitate the annotation process, including 1) an interactive annotation tool that defines the annotation format and automatically unifies all annotations as JSON files (Please see §B.1 for our annotation format), 2) a GitHub repository to coordinate the task submission, reviewing, and discussion process, which was inspired by BIG-bench (Srivastava et al., 2022), and 3) a visualization tool that allows annotators to browse the existing tasks and the evaluation results of representative vision language models (VLMs). We coordinate all the annotators to ensure they understand our expectations and continuously improve our tools.
@@ 3.2 Metrics for answers in diverse output formats
Rule-based metrics. When there is a unique answer under the question context or the correctness of the answer can be verified by rules (e.g., if the generated story/poetry meets the desired formats, or if the generated code can pass the given test cases), we implement rule-based metrics for evaluation. To satisfy the needs of all tasks submitted by our annotators, we end up with a suite of over 40 rule-based metrics. Robust string parsing is also implemented to extract the answer from the model’s response. We conduct a sanity check to ensure the correct implementation of all rule-based metrics. Specifically, we create an “oracle” model that always returns the ground truth annotation. We then compute its score over all tasks evaluated with rule-based metrics. The sanity check is passed when the “oracle” model gets a 1.0 score. See §D.4 for details.
LLM-assisted metrics. For open-ended tasks that do not have a unique answer, we instead employ an LLM-assisted metric (Zheng et al., 2023; Li et al., 2023a). We design a per-task evaluation prompt template and fill in the tailored evaluation criteria for each task. The LLM is instructed to compare the model response with the reference answer and assign a score from 1 to 10. The score is then normalized into [0, 1] to be consistent with the other metrics. See §D.3 for details.
@@ 3.3 Multi-dimensional Keywords for fine-grained analysis
MVBench (Li et al., 2024e) Repurposed Existing Video Multi-choice (MC) 1 20
@@ 3.4 Dataset statistics and comparison with other benchmarks
MEGA-Bench contains 505 real-world tasks with 8,186 manually annotated or repurposed samples. Even for repurposed data, considerable effort is needed to convert the original annotations into specific task descriptions, diverse output formats, and additional instructions to include auxiliary information and about the formatting. Figure 2 shows the task distribution of all five dimensions, and the detailed breakdown statistics are in §C.2.
@@ 4 Experiments
We evaluate 22 VLMs with multi-image support on MEGA-Bench. §4.1 describes the models being assessed and the evaluation pipeline. §4.2 presents the quantitative evaluation results with a fine-grained analytical breakdown. §4.3 discusses the impact of the number of examples per task on the evaluation results. §4.4 provides error analysis based on GPT-4o. Detailed inspections of model behaviors with concrete error examples can be found in §F of the Appendix.
@@ 4.1 Evaluation Settings
Evaluated models. We evaluate a diverse range of large multimodal models. The proprietary models assessed include GPT-4o (0513) and GPT-4o mini (OpenAI, 2024a), Claude-3.5-Sonnet (0620 and 1022) (Anthropic, 2024a; b), Gemini-1.5-Pro (002) and Gemini-1.5-Flash (002) (Google, 2024a). For open-source models, we mainly focus on large flagship ($>$70B parameters) and small-to-medium efficiency models. The large models include Qwen2-VL-72B (Alibaba, 2024), InternVL2-Llama3-76B (Chen et al., 2024d), LLaVA-OneVision-72B (Li et al., 2024a), and NVLM (Dai et al., 2024). The medium-scale models comprise Qwen2-VL-(7B/2B) (Alibaba, 2024), Pixtral 12B (Mistral, 2024), Aria (Li et al., 2024d), InternVL2-(8B/2B) (Chen et al., 2024d), Phi-3.5-Vision (Abdin et al., 2024), MiniCPM-V2.6 (Yao et al., 2024), LLaVA-OneVision-7B (Li et al., 2024a), Llama-3.2-11B Meta (2024), Idefics3-8B-Llama3 (Laurençon et al., 2024), and Aquila-VL-2B-llava-qwen (Gu et al., 2024).
Evaluation pipeline. MEGA-Bench has diverse and flexible formats. To ensure the models have clear instructions on the output format, we provide all evaluated VLMs with a one-shot in-context example. For each query, we fill in a pre-defined prompt template with the task instructions written by the task annotators, the 1-shot example, and the concrete query question. Since this one-shot example’s primary purpose is to illustrate the output format, we allocate only a tiny portion of the total image budget for it. For each model, we conduct experiments with and without Chain-of-Thought (CoT) prompting (Wei et al., 2022) for the Core tasks (the one-shot example of Open-ended tasks already contains CoT demonstrations). The prompt templates and other evaluation details (e.g., the frame sampling strategy for video inputs) are in §D.
Single-image setting. Our evaluation pipeline focuses on models with multi-image support. To properly evaluate models trained mainly for single-image use cases, we create a single-image setting using the single-image tasks of MEGA-Bench. See §A for the detailed results and analyses.
Table 2: The main results of different models on the Core and Open-ended subset of MEGA-Bench, with 440 and 65 tasks, respectively. We report the macro mean scores across all tasks in each set. The overall score is the weighted average of the Core and Open-ended scores. When computing the overall score, we use the higher Core score from ‘w/o CoT’ and ‘w/ CoT’. Some open-source models perform more reasonably when the input contains only a single image. Please see §A for results and analyses of our single-image setting.
Model Eval Tier Open Source Core (rule eval) Open-ended (GPT eval) Overall
@@
Claude-3.5-Sonnet (1022) (Anthropic, 2024b) Flagship No 49.20 52.59 65.63 54.27
GPT-4o (0513) (OpenAI, 2024a) Flagship No 52.03 52.65 64.78 54.21
Claude-3.5-Sonnet (0620) (Anthropic, 2024a) Flagship No 48.80 50.41 63.74 52.13
Gemini-1.5-Pro-002 (Google, 2024b) Flagship No 46.99 48.22 58.58 49.55
Gemini-1.5-Flash-002 (Google, 2024b) Efficiency No 41.90 41.89 56.91 43.82
GPT-4o mini (OpenAI, 2024b) Efficiency No 39.85 40.77 58.65 43.07
Qwen2-VL-72B (Alibaba, 2024) Flagship Yes 46.41 45.42 56.40 47.70
InternVL2-Llama3-76B (Chen et al., 2024d) Flagship Yes 35.02 35.63 51.93 37.73
LLaVA-OneVision-72B (Li et al., 2024a) Flagship Yes 31.99 29.74 45.99 33.79
NVLM-72B (Dai et al., 2024) Flagship Yes 24.21 21.59 34.78 25.57
Qwen2-VL-7B (Alibaba, 2024) Efficiency Yes 34.80 32.93 43.96 35.98
Pixtral-12B (Mistral, 2024) Efficiency Yes 31.91 31.36 45.66 33.68
Aria-MoE-25B (Li et al., 2024d) Efficiency Yes 30.49 28.90 51.03 33.13
InternVL2-8B (Chen et al., 2024d) Efficiency Yes 25.96 24.09 39.79 27.74
Phi-3.5-Vision-4B (Abdin et al., 2024) Efficiency Yes 23.27 23.00 39.48 25.36
MiniCPM-V2.6-8B (Yao et al., 2024) Efficiency Yes 22.88 22.96 41.73 25.38
LLaVA-OneVision-7B (Li et al., 2024a) Efficiency Yes 22.41 21.36 33.98 23.90
Qwen2-VL-2B (Alibaba, 2024) Efficiency Yes 16.45 20.88 31.54 22.25
Llama-3.2-11B (Meta, 2024) Efficiency Yes 10.04 16.00 31.73 18.02
Aquila-VL-2B-llava-qwen (Gu et al., 2024) Efficiency Yes 16.32 16.00 24.57 17.38
InternVL2-2B (Chen et al., 2024d) Efficiency Yes 9.09 13.14 23.86 14.52
@@ 4.2 Main Results with breakdown analysis
Flagship models. Unlike the results on recent benchmarks like MMMU-Pro (Yue et al., 2024b) where GPT-4o (0513) and Claude-3.5-Sonnet (0620) get close scores, GPT-4o (0513) outperforms Claude-3.5-Sonnet (0620) with a clear margin on MEGA-Bench ($>$ 2%). Investigating the breakdown results, we observe that GPT-4o (0513) wins in most applications/skills except for coding, math, and planning-related tasks, where the answers are typically in a “structured” output format, as shown in Figure 5. The recent update of Claude-3.5-Sonnet (1022) makes consistent improvements across almost all dimensions, especially in planning-related tasks and those with infographics/UI/photographs inputs, and slightly surpasses GPT-4o (0513) in the overall score ($<$ 0.1%). The “Planning” application keyword contains tasks like symbolic planning (Zhu et al., 2024), navigation (Ku et al., 2020), chess games (Fu et al., 2024b), puzzle games (e.g., maze, Sudoku, etc.), and even the best models get low scores.
One typical observation of Claude-3.5-Sonnet models is that they tend to be meticulous and refuse to answer routine knowledge or commonsense questions, such as the name and nationality of famous actors. The bottom radar maps show that they fall behind in knowledge, information extraction, and commonsense reasoning compared to GPT-4o, partially because of this refusal behavior.
@@
Efficiency models. Figure 6 analyzes the results on efficiency models. In general, Gemini-1.5-flash-002 has the best performance with exceptional scores in Science and Metrics applications. The Metrics keyword contains tasks such as rating the quality of GenAI results (He et al., 2024; Jiang et al., 2024b) and requires deep multimodal reasoning and commonsense. However, its performance on UI-related inputs and information extraction tasks falls behind GPT-4o mini.
Chain-of-Thought. An interesting finding is that the CoT prompt (See §D) effectively guides all proprietary models to generate a detailed reasoning process, and flagship-tier proprietary models all obtain better performance on the Core set. However, it has almost no effect on most open-source models. For example, the Qwen2-VL, InternVL2, and LLaVA-OneVision models rarely produce reasoning when given a CoT instruction, and sometimes get confused about the required format after generating the reasoning process, leading to a lower score on the Core set.
Some open-source models get comparatively low scores for their parameter count. Llama-3.2-11B has difficulty leveraging the one-shot example to understand the correct output format and tends to generate a long descriptive sentence instead. This issue is alleviated under the CoT setting because the prompt provides extra instructions on the output format beyond the one-shot example, requesting the model to strictly separate the reasoning process from the final answer. Idefics3 frequently repeats the example answer from the one-shot demonstration. We suspect the reason for this problem is the poor support for multi-image (as our query contains at least two images, including the one-shot example) since it can generate reasonable responses with a single-image input. §A in the Appendix presents a single-image setting of MEGA-Bench and conducts further analyses.
@@ 4.3 Analysis on the number of samples per task
As discussed in §1, one of MEGA-Bench’s goals is to optimize the inference cost while still producing detailed multi-dimensional breakdown analysis. Therefore, we prioritize expanding the number of tasks over adding many examples per task in the benchmark construction process. To understand the robustness of the benchmark score with around 15 examples per task, we obtained bootstrap distributions (Efron & Tibshirani, 1994; Hesterberg, 2011) of the model scores for our Core set with the CoT prompting. We did this by taking a random subset of the model’s responses of size $n$ ($n=3,5,\dots,13,15$) with replacement for each task and calculating the task-level macro-mean scores. To ensure the bootstrap distribution was numerically stable, we ran 10,000 Monte Carlo simulations. Figure 7 (left) shows that the variance in model scores rapidly narrows as the number of examples per task increases. As the number of examples per task increases beyond 7, the marginal return in variance reduction diminishes.
Figure 7: (Left) The bootstrap distribution of benchmark scores of a subset of models as we gradually increase the bootstrap sample size of the number of examples per task. We use the results with the CoT prompting. Claude 3.5 Sonnet (1022) and Gemini 1.5 Pro (002) are used in this figure. (Right) The task-wise error distribution of GPT-4o (0513) over a subset of 255 Core tasks.